# Strategic Classification with Non-Linear Classifiers

### v1

@@ B.3 Proof of Proposition 1
Let $x$ be a point on the decision boundary of $h$ with signed curvature $\kappa\geq-1/\alpha$. The effective curvature of the corresponding ${\hat{x}}$ on the boundary of $h_{\Delta}$ is given by:
@@
While an equivalent form of Prop. 1 has already been proven in the field offset curves [12], we nonetheless reprove it here to offer some intuition behind it. In the normal one-to-one mapping, because both the center of curvature and $\nabla_{h}(x)$ lie on the normal line to $h$ at $x$, strategic movement can be seen as increasing the signed radius of curvature $r_{osc}$ by $\alpha$ (where we will define the radius of curvature as negative if the curvature is negative at $x$). In higher dimensions, this is equivalent to increasing each of the radii of curvature by $\alpha$ (see Appendix A.1). We therefore have:
@@
Note that Eq. 3 has a horizontal asymptote at $\kappa_{\Delta}=1/\alpha$ because the effective curvature cannot exceed 1/$\alpha$ (see Sec. 5.3). Additionally, Eq. 3 has a vertical asymptote at $\kappa=-1/\alpha$ because any point with lower curvature has no effect on $h_{\Delta}$ (see Appendix B.2). When $\kappa=-1/\alpha$, $x$ is mapped to a non-smooth kink of infinite curvature. ∎
@@ B.4 Proof of Proposition 2
Let $g$ be a candidate classifier. If there exists a point $x$ on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}(x)$ are reachable from some point in ${\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
@@
In order for $g$ to be the effective classifier of regular classifier $h$, we must have:
@@
    For each point $z$ that is positively classified by $g$, there must be a point $z^{\prime}\in B_{\alpha}(z)$ which is classified as positive by $h$. In other words, $z$ must be able to acquire a positive label by staying in place or strategically moving within $B_{\alpha}(z)$.
@@
    For each point $z$ that is negatively classified by $g$, there cannot be any points $z^{\prime}\in B_{\alpha}(z)$ which are classified as positive by $h$. In other words, $z$ must not be able to acquire a positive label by staying in place or strategically moving within $B_{\alpha}(z)$.
Assume there exists a point $x$ on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}(x)$ are reachable from ${\cal{X}}_{0}(g)$ and that there exists $h$ such that $g=h_{\Delta}$. Since $x$ is on the decision boundary of $g$, there exists a point $x+\delta,\delta\to 0$ which is negatively classified by $g$. For requirement 1 and 2 to hold, there must be a point $x^{\prime}\in S=\{z:z\in B_{\alpha}(x),z\not\in B_{\alpha}(x+\delta)\}$ that is positively classified by $h$. Since $S\subset S_{\alpha}(x)$, we can expand the claim to say that there must be a point $x^{\prime}\in S_{\alpha}(x)$ that is positively classified by $h$. However, if all points $x^{\prime}\in S_{\alpha}(x)$ are reachable by some point $z$ that is negatively classified by $g$, then none of the points $x^{\prime}\in S_{\alpha}(x)$ can be positively classified by $h$ (by requirement 2). Therefore, if there exists a point $x$ on the decision boundary of $g$ for which every point on the sphere $S_{\alpha}(x)$ is reachable by some point that is negatively classified by $g$, then $g$ cannot be a possible effective classifier. ∎
@@ B.5 Proof of Proposition 3
Let $g$ be a candidate classifier. If there exists a point $x$ on the decision boundary of $g$ where (i) $g$ is smooth, and (ii) the point ${\hat{x}}=x+\alpha\hat{n}_{x}$ is reachable by some $x^{\prime}\in{\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
@@
By proposition 2, for each point $x$ on the decision boundary of $g$, there must be a point $x^{\prime}$ on $S_{\alpha}(x)$ which is classified positively by $h$ and also not reachable by any point negatively classified by $g$. As a result, the ball $B_{\alpha}(x^{\prime})$, which is the set of points that can move to $x^{\prime}$, must not intersect the decision boundary and must furthermore be strictly tangent to the decision boundary at $x$. Given that the decision boundary is smooth at $x$, this means that the normal to the curve at $x$ must run along the radius of $B_{\alpha}(x^{\prime})$ from $x$ to $x^{\prime}$, implying that $x^{\prime}=x\pm\alpha\hat{n}_{x}$. Because the strategic movement is in the direction of a positive classification, the only possible option for $x$ is to move to $x^{\prime}=x+\alpha\hat{n}_{x}$. However, if $x^{\prime}=x+\alpha\hat{n}_{x}$ is reachable by some point that is negatively classified by $h_{\Delta}$, then the associated $h$ would not have classified $x^{\prime}$ as positive, so $x$ would not have been a positively classified point by $h_{\Delta}$. ∎
@@ B.6 Effective Class of Polytope Family
Figure 7: (Left) Each face of a negative-inside polytope is mapped to a new linear face, though the resulting $h_{\Delta}$ may be a polytope with fewer faces and vertices due to wipeout. (Right) Each face of a positive-inside polytope is also mapped to a new linear face, though each vertex will have to expand to fill the gap between theses faces.
Each polytope is simply a combination of $n$ linear faces. As such, when a polytope has a negative inside, each face will be mapped to a new linear face (with some wipeout of part or all of the face) leading to an $h_{\Delta}$ that is a smaller-radius polytope, albeit with potentially fewer faces and edges (see Fig. 7 (Left)). Although the faces of a positive-inside polytope also map to new linear faces, each vertex will have to expand to fill the gap between theses new faces. As such, $h_{\Delta}$ is now a larger polytope whose corners are partial to a ball (see Fig. 7 (Right)). For the $\ell_{1}$ and $\ell_{\infty}$ costs, $h_{\Delta}$ therefore remains a polytope since the $\ell_{1}$ and $\ell_{\infty}$ balls are peicewise-linear themselves. For the $\ell_{2}$ cost, $h_{\Delta}$ is now a polytope with rounded corners. In either case, the decision boundary of each $p^{k}_{\Delta}\in P^{k}_{\Delta}$ can be seen as the intersection of a radius-scaled $p^{k}$ and $k$ balls centered at each vertex of $p^{k}$.
@@ B.7 Proof of Proposition 4
For each class $Q^{k}_{s}$ of negative-inside $k$-vertex polytopes bounded by a radius of $s$, $\mathrm{VC}(Q^{k}_{s})\geq\mathrm{VC}(Q_{s,\Delta}^{k})$. Moreover, if $s=\infty$, then $\mathrm{VC}(Q^{k}_{s})=\mathrm{VC}(Q_{s,\Delta}^{k})$ as implied by Thm. 2.
@@ B.8 Proof of Theorem 1
Figure 8: Example $h$ from a class $H$ where $\mathrm{VC}(H)=\infty$, but $\mathrm{VC}(H_{\Delta})=0$. The grey dashed curve is the $h_{\Delta}$ of any $h$ like the one above with the same small gap where the decision boundary remains above the dotted purple line.
@@
We provide two examples. For the first example, fix $\alpha<1$ and consider the class of $\mathbb{R}^{2}$ classifiers $H=\{h_{S}\mid S\subseteq\mathbb{Z}^{2}\}$ where $h_{S}=\mathds{1}{\{{x\not\in B_{\alpha-\delta}(z)\,\,\forall z\in S}\}}$. Each classifier in this model class contains small, non-overlapping negative regions centered at lattice points. Thus, the negative regions of all $h\in H$ will disappear under strategic movement leaving $h_{\Delta}(x)=1\,\,\forall h\in H$. Despite the fact that $H$ shatters $\mathbb{Z}^{2}$ (and therefore has $\mathrm{VC}(H)=\infty$), $H_{\Delta}$ cannot shatter a single point (and therefore has $\mathrm{VC}(H_{\Delta})=0$).
Another example where $h_{\Delta}$ is perhaps more natural can be seen in Fig. 8. In this case, $H$ is the class of classifiers whose decision boundaries can be written as
$h_{i}\big{(}(x_{1},x_{2})\big{)}=\begin{cases}0&x_{1}\leq-k,\\ f_{i}\big{(}(x_{1},x_{2})\big{)}&-k\leq x_{1}\leq k\\ 0&x_{1}\geq k\end{cases}$ (4)
where $k$ is some constant less than $\alpha$ and each function $f_{i}\big{(}(x_{1},x_{2})\big{)}$ satisfies $f_{i}\big{(}(x_{1},x_{2})\big{)}\geq\alpha^{2}-x_{1}^{2}-x_{2}^{2}$ for all $-k\leq x_{1}\leq k$. In this case, decision boundary points in the gap will get wiped out on all $h\in H$ leaving the exact same $h_{\Delta}$. Because $H$ is nearly unrestricted in the range $-k\leq x_{1}\leq k$, it is capable of shattering infinite points that lie in this range and has infinite VC dimension. On the other hand, because $|H_{\Delta}|=1$, we get $\mathrm{VC}(H_{\Delta})=0$. ∎
@@ B.9 Proof of Theorem 2
Note that the family of polynomials and polytopes are clearly closed to input scaling. Furthermore, most classes of neural networks are as well since $h(ax)$ can be realized by scaling the first weight vector by $a$.
@@
Intuitively, the reason that the VC dimension cannot decrease is that for any $H$ with VC dimension $k$ that is closed to input scaling, we can find a set $S$ of $k$ point that can be shattered by $H$ where none of the points $s\in S$ strategically move under any of the $h\in H$ used to shatter $S$. As a result, $H_{\Delta}$ will also shatter $S$ so its VC dimension is at least $k$.
@@
Since both $S$ and $G$ are finite sets and negative points face a strictly positive cost to obtain a positive classification, we get that $c_{neg}(S,G)>0$ Additionally, because the VC dimension is indifferent to input scaling [^6] Input scaling only changes the distance between points, but does not change their relative positions, so the class $H$ and its input scaled class $H^{\prime}$ will have the same VC dimension., if we scale both the classifiers and the points, then $G^{\prime}=\{g(ax):g(x)\in G\}$ will shatter $S^{\prime}=\{ax:x\in S\}$ with $c_{neg}(S^{\prime},G^{\prime})=a\cdot c_{neg}(S,G)$. Note that because $H$ is closed to input scaling, $G^{\prime}\subset H$ as well. Furthermore, if we choose $a=\frac{\alpha+1}{c_{neg}(S,G)}$, then $c_{neg}(S^{\prime},G^{\prime})=\alpha+1$. Note that if $c_{neg}(S^{\prime},G^{\prime})>\alpha$, no point $s^{\prime}\in S^{\prime}$ will move strategically under any $g^{\prime}\in G^{\prime}$, meaning that $g^{\prime}_{\Delta}(s^{\prime})=g^{\prime}(s^{\prime})\,\,\forall g^{\prime}\in G,\,\,s^{\prime}\in S^{\prime}$. As a result, $G^{\prime}_{\Delta}$ will also shatter $S^{\prime}$ so $\mathrm{VC}(G^{\prime}_{\Delta})\geq|S^{\prime}|=k$. Since $G^{\prime}_{\Delta}\subset H_{\Delta}$, we get that
$\mathrm{VC}(H_{\Delta})\geq\mathrm{VC}(G^{\prime}_{\Delta})\geq k=\mathrm{VC}(H)$ (6)
@@ B.10 Proof of Theorem 3
Consider either the $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ costs and let $H\subseteq P^{k}$, where $P^{k}$ is the set of all $k$-vertex polytopes in $R^{d}$. Then $\mathrm{VC}(H_{\Delta})=O(kd^{2}\log k)$.
@@
Let classes $C_{1},C_{2},..C_{k}$ have VC dimensions $V_{1},V_{2},...V_{k}$ and $V=\sum_{i=1}^{k}V_{i}$. Let $C_{U}=\bigcup_{i}\,C_{i}$ and Let $C_{I}=\bigcap_{i}\,C_{i}$. Then both $C_{U}$ and $C_{I}$ have VC dimension upper bounded by $O\big{(}V\cdot log(k)\big{)}$ [32].
We note that by combining Lemmas 2 and 3, the class $H^{Ball}_{k}$ of the union of $k$ $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ balls has a VC dimension of $O\big{(}k\,\,d\,\,log(k)\big{)}$. As seen in Appendix B.6, the effective classifier of a k-vertex positive-inside polytope is the intersection of a larger radius k-vertex positive-inside polytope and $k$ $\ell_{p}$ balls centered at each of the vertices. Therefore, $H_{\Delta}\subset P^{k}\,\cap H^{ball}_{k}$ so :

### v2

@@ B.3 Proof of Proposition 1
Figure 6: (Top) Graph of Eq. 3 with asymptotes represented as dashed lines. (Bottom) Depictions of curvature mappings for increasing $\kappa$.
@@
Let z{z} be a point on the decision boundary of $h$ with signed curvature $\kappa\geq-1/\alpha$. The effective curvature of the corresponding x{x} on the boundary of $h_{\Delta}$ is given by:
@@
While an equivalent form of Prop. 1 has already been proven in the field offset curves [12], we nonetheless reprove it here to offer some intuition behind it. In the normal, one-to-one mapping, because both the center of curvature and $\nabla_{h}({z})$ lie on the normal line to $h$ at z{z} (see Appendix B.1), strategic movement can be seen as increasing the signed radius of curvature, $r_{osc}$, by $\alpha$ (where we will define the radius of curvature as negative if the curvature is negative at z{z}). In higher dimensions, this is equivalent to increasing each of the radii of curvature by $\alpha$ (see Appendix A.1). We therefore have:
@@
[…]
@@ B.4 Proof of Proposition 2
Let $g$ be a candidate classifier. If there exists a point x{x} on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}({x})$ are reachable from some point in ${\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
@@
In order for $g$ to be the effective classifier of the regular classifier $h$, we must have:
@@
    For each point q{q} that is positively classified by $g$, there must be a point ${q}^{\prime}\in B_{\alpha}({q})$ which is classified as positive by $h$. In other words, q{q} must be able to acquire a positive label by staying in place or strategically moving within $B_{\alpha}({q})$.
@@
    For each point q{q} that is negatively classified by $g$, there cannot be any points ${q}^{\prime}\in B_{\alpha}({q})$ which are classified as positive by $h$. In other words, q{q} must not be able to acquire a positive label by staying in place or strategically moving within $B_{\alpha}({q})$.
Assume there exists a point x{x} on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}({x})$ are reachable from ${\cal{X}}_{0}(g)$ and that there exists $h$ such that $g=h_{\Delta}$. Since x{x} is on the decision boundary of $g$, there exists a point ${x}+\delta,\delta\to 0$ that is negatively classified by $g$. For requirements 1 and 2 to hold, there must be a point $x^{\prime}\in S=\{{q}:{q}\in B_{\alpha}({x}),{q}\not\in B_{\alpha}({x}+\delta)\}$ that is positively classified by $h$. Since $S\subset S_{\alpha}({x})$, we can expand the claim to say that there must be a point $x^{\prime}\in S_{\alpha}({x})$ that is positively classified by $h$. However, if all points $x^{\prime}\in S_{\alpha}({x})$ are reachable by some point q{q} that is negatively classified by $g$, then none of the points $x^{\prime}\in S_{\alpha}({x})$ can be positively classified by $h$ (by requirement 2). Therefore, if there exists a point x{x} on the decision boundary of $g$ for which every point on the sphere $S_{\alpha}({x})$ is reachable by some point that is negatively classified by $g$, then $g$ cannot be a possible effective classifier. ∎
@@ B.5 Proof of Proposition 3
Let $g$ be a candidate classifier. If there exists a point x{x} on the decision boundary of $g$ where (i) $g$ is smooth, and (ii) the offset point ${\hat{x}}={x}+\alpha\hat{n}_{x}$ is reachable by some other $x^{\prime}\in{\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
@@
By Prop. 2, for each point x{x} on the decision boundary of $g$, there must be a point $x^{\prime}$ on $S_{\alpha}({x})$ that is classified positively by $h$ and also not reachable by any point negatively classified by $g$. As a result, the ball $B_{\alpha}(x^{\prime})$, which is the set of points that can move to $x^{\prime}$, must not intersect the decision boundary and must furthermore be strictly tangent to the decision boundary at x{x}. Given that the decision boundary is smooth at x{x}, this means that the normal to the curve at x{x} must run along the radius of $B_{\alpha}(x^{\prime})$ from x{x} to $x^{\prime}$, implying that $x^{\prime}={x}\pm\alpha\hat{n}_{x}$. Because the strategic movement is in the direction of a positive classification, the only possible option is for x{x} to move to $x^{\prime}={x}+\alpha\hat{n}_{x}$. However, if $x^{\prime}={x}+\alpha\hat{n}_{x}$ is reachable by some point that is negatively classified by $h_{\Delta}$, then the associated $h$ would not have classified $x^{\prime}$ as positive, so x{x} would not have been a positively classified point by $h_{\Delta}$. ∎
@@ B.6 Effective Class of Polytope Family
Figure 7: (Left) Each face of a negative-inside polytope is mapped to a new linear face, though the resulting $h_{\Delta}$ may be a polytope with fewer faces and vertices due to wipeout. (Right) Each face of a positive-inside polytope is also mapped to a new linear face, though each vertex will have to expand to fill the gap between these faces.
Each polytope is simply a combination of $n$ linear faces. As such, when a polytope has a negative inside, each face will be mapped to a new linear face (with some wipeout of part or all of the face) leading to an $h_{\Delta}$ that is a smaller-radius polytope, albeit with potentially fewer faces and edges (see Fig. 7 (Left)). Although the faces of a positive-inside polytope also map to new linear faces, each vertex will have to expand to fill the gap between these new faces. As such, $h_{\Delta}$ is now a larger polytope whose corners are partial to a ball (see Fig. 7 (Right)). For the $\ell_{1}$ and $\ell_{\infty}$ costs, $h_{\Delta}$ therefore remains a polytope since the $\ell_{1}$ and $\ell_{\infty}$ balls are piecewise-linear themselves. For the $\ell_{2}$ cost, $h_{\Delta}$ is now a polytope with rounded corners. In either case, the decision boundary of each $p^{k}_{\Delta}\in P^{k}_{\Delta}$ can be seen as the intersection of a radius-scaled $p^{k}$ and $k$ balls centered at each vertex of $p^{k}$.
@@ B.7 Proof of Proposition 4
Let $Q^{k}_{s}$ be the class of negative-inside $k$-vertex polytopes bounded by a radius of $s$, then $\mathrm{VC}(Q^{k}_{s})\geq\mathrm{VC}(Q_{s,\Delta}^{k})$. Moreover, if $s=\infty$, then $\mathrm{VC}(Q^{k}_{s})=\mathrm{VC}(Q_{s,\Delta}^{k})$ as implied by Thm. 2.
@@ B.8 Proof of Theorem 1
Figure 8: Example $h$ from a class $H$ where $\mathrm{VC}(H)=\infty$, but $\mathrm{VC}(H_{\Delta})=0$. The gray dashed curve is the $h_{\Delta}$ of any $h$ like the one above. Each such $h$ has the same small gap where the decision boundary remains above the dotted purple curve.
@@
We provide two examples. For the first example, fix $\alpha<1$ and consider the class of $\mathbb{R}^{2}$ classifiers $H=\{h_{S}:S\subseteq\mathbb{Z}^{2}\}$ where $h_{S}=\mathds{1}{\{{x\not\in B_{\alpha-\delta}(s)\,\,\forall s\in S}\}}$. Each classifier in this model class contains small, non-overlapping negative regions centered at lattice points. Thus, the negative regions of all $h\in H$ will disappear under strategic movement, leaving $h_{\Delta}(x)=1\,\,\forall h\in H$. Despite the fact that $H$ shatters $\mathbb{Z}^{2}$ (and therefore has $\mathrm{VC}(H)=\infty$), $H_{\Delta}$ cannot shatter a single point (and therefore has $\mathrm{VC}(H_{\Delta})=0$).
Another example where $h_{\Delta}$ is perhaps more natural can be seen in Fig. 8. In this case, $H$ is the class of classifiers with narrow gaps whose decision boundaries can be written as
[…]
where $k$ is some constant less than $\alpha$ and each function $f_{i}(x_{1})$ satisfies $f_{i}(x_{1})\geq\sqrt{\alpha^{2}-x_{1}^{2}}+\sqrt{\alpha^{2}-k^{2}}$ for all $k\leq x_{1}\leq k$ (i.e., $f_{i}(x)$ lies above the radius-$\alpha$ arc between (k,0)(-k,0) and $(k,0)$). In this case, the decision boundary points in the gap will get wiped out on all $h\in H$, leaving the exact same $h_{\Delta}$. Because $H$ is nearly unrestricted in the range $k\leq x_{1}\leq k$, it is capable of shattering infinite points that lie in this range and has infinite VC dimension. On the other hand, because $|H_{\Delta}|=1$, we get $\mathrm{VC}(H_{\Delta})=0$. ∎
@@ B.9 Proof of Theorem 2
Note that the families of polynomials and polytopes are clearly closed to input scaling. Furthermore, most classes of neural networks are as well since $h(ax)$ can be realized by scaling the first weight vector by $a$.
@@
Intuitively, the reason that the VC dimension cannot decrease is that for any $H$ with a VC of dimension $k$ that is closed to input scaling, we can find a set $S$ of $k$ points that can be shattered by $H$, where none of the points $s\in S$ strategically move under any of the $h\in H$ used to shatter $S$. As a result, $H_{\Delta}$ will also shatter $S$, so its VC dimension is at least $k$.
@@
Since both $S$ and $G$ are finite sets and negative points face a strictly positive cost to obtain a positive classification, we get that there exists a unique value $c_{neg}(S,G)>0$ . Additionally, because the VC dimension is indifferent to input scaling [^5] Input scaling only changes the distance between points, but does not change their relative positions, so the class $H$ and its input-scaled class $H^{\prime}$ will have the same VC dimension., if we scale both the classifiers and the points, then $G^{\prime}=\{g(ax):g(x)\in G\}$ will shatter $S^{\prime}=\{ax:x\in S\}$ with $c_{neg}(S^{\prime},G^{\prime})=a\cdot c_{neg}(S,G)$. Note that because $H$ is closed to input scaling, $G^{\prime}\subset H$ as well. Furthermore, if we choose $a=\frac{\alpha+1}{c_{neg}(S,G)}$, then $c_{neg}(S^{\prime},G^{\prime})=\alpha+1$. Note that if $c_{neg}(S^{\prime},G^{\prime})>\alpha$, no point $s^{\prime}\in S^{\prime}$ will move strategically under any $g^{\prime}\in G^{\prime}$, meaning that $g^{\prime}_{\Delta}(s^{\prime})=g^{\prime}(s^{\prime})\,\,\forall g^{\prime}\in G,\,\,s^{\prime}\in S^{\prime}$. As a result, $G^{\prime}_{\Delta}$ will also shatter $S^{\prime}$ so $\mathrm{VC}(G^{\prime}_{\Delta})\geq|S^{\prime}|=k$. Since $G^{\prime}_{\Delta}\subset H_{\Delta}$, we get that
$\mathrm{VC}(H_{\Delta})\geq\mathrm{VC}(G^{\prime}_{\Delta})\geq k=\mathrm{VC}(H)$
@@ B.10 Proof of Theorem 3
Let $H^{m,k}$ be a class of piecewise-linear classifiers with at most m segments and k intersections of linear segments. Then $\mathrm{VC}(H_{\Delta}^{m,k})=O\big(d\,m\,log(m)+\nu_{p}\,k\,log(k)\big)$, where $\nu_{p}$ is the VC of the class of $\ell_{p}$ norm balls, and is $\Theta(d)$ for $p=1,2,\infty$.
@@
[…]
We note that by combining Lemmas 2 and 3, the class $H^{Ball}_{k}$ of the union of up to $k$ $\ell_{p}$ balls has a VC dimension of $O\big(\nu_{p}\,\,k\,\,log(k)\big)$, which becomes $O\big(d\,\,k\,\,log(k)\big)$ in the case of either the $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ norm cost functions. Additionally, the class $H^{Lin}_{m}$ of piecewise linear models with up to $m$ segments has a VC dimension of $O\big(d\,\,m\,\,log(m)\big)$.  [^6] Note that combining the two technically proves the bounds for the classes of the union of exactly $k$ balls and exactly $m$ segments, respectively, though we can easily expand them to the classes of up to $k$ balls and up to $m$ segments by recognizing that some of the balls/segments from classes $C_{1},C_{2},...C_{k}$ may be the exact same.
Because linear classifiers map to new shifted linear classifiers, all we have to do to determine the effective classifier of a piecewise-linear $h$ is to combine (take the union of) the effective segments of each linear segment in $h$ and the potential $\ell_{p}$ ball artifacts around each intersection. Note that, as seen in Fig. 7 (Left), some of the effective segments may be wiped out in this process, and, as seen in Fig. 7 (Right), only intersections that open to the positive region will induce $\ell_{p}$ ball artifacts in $h_{\Delta}$. Therefore, the $h_{\Delta}$ of a piecewise linear $h$ with $m$ segments and $k$ intersections will be the intersection of a piecewise linear classifier $g$ with up to $m$ segments and up to $k$ $\ell_{p}$ ball artifacts. As a result, we can write $H_{\Delta}\subseteq G\,\cap\,H^{Ball}_{k}$, where $G\subseteq H^{Lin}_{m}$. Therefore:
$\mathrm{VC}(H_{\Delta})$ $\leq\mathrm{VC}\big(G\,\cap H^{ball}_{k}\big)$
$=O\big(\mathrm{VC}(G)\,+\mathrm{VC}(H^{ball}_{k})\big)$
$\leq O\big(\mathrm{VC}(H^{Lin}_{m})\,+\mathrm{VC}(H^{ball}_{k})\big)$
$=O\big(d\,m\,log(m)+\nu_{p}\,\,k\,\,log(k)\big)$
Consequently, if $H$ is learnable, then $H_{\Delta}$ is too.
B.11 Proof of Theorem 4
Theorem 4:
[…]
Proof.
As seen in Appendix B.6, the effective classifier of a k-vertex polytope is the intersection of a larger radius (or smaller for negative inside polytopes) k-vertex polytope and $k$ $\ell_{p}$ balls centered at each of the vertices. Therefore, $H_{\Delta}\subset P^{k}\,\cap H^{ball}_{k}$, so by Lemmas 2 and 3 (and the VC dimension of $H^{ball}_{k}$ derived in Appendix B.10):